Random variables X, ~ e.rp(λ.) are independent of each other, where i=1,2,3 and that lambda _{i}>0,orall i

Calculate P(X1<X2<X3)
Answer #1

We know that,

Min(X1, X2,, ...., Xn) ~ Exp(lambda_1 + lambda_2 + ...+ lambda_n)

and,

P(Xi = Min(X1, X2, ...., Xn)) = lambda_i / (lambda_1 + lambda_2 + ...+ lambda_n)

and

P(Xi < Xj) = lambda_i / (lambda_i + lambda_j)

Now,

P(X1<X2<X3) = P(X1 = min(X1,X2,X3))· P(X2 < X3|X1 = min(X1,X2,X3))

By memoryless property of exponential distribution,

P(X2 < X3|X1 = min(X1,X2,X3) = P(X2 < X3)

Thus,

P(X1<X2<X3) = P(X1 = min(X1,X2,X3))· P(X2 < X3)